dina and her family are going camping.The campsite is 3/4 miles away.They walk at a stedy speed of 2 1/2 mi/h.How many minutes would it take to get to the campsite?
1
Expert's answer
2011-10-13T12:55:40-0400
S=3/4 mi
V=2*1/2*mi/h
t-?

t=S/V
t=3/4*5/2=3/10 h
t=3*60/10=18 min